OpenStudy (anonymous):

Verify the following identity: sin(theta)cos^2(theta)+sin^3(theta)=sin(theta) Please include steps! :/ Thank you!

OpenStudy (amistre64):

split it all into sins

OpenStudy (amistre64):

cos^2 = 1 - sin^2 right?

OpenStudy (anonymous):

rightt

OpenStudy (amistre64):

then distribute the sin thru and cancel the ^3s

OpenStudy (amistre64):

make sense?

OpenStudy (anonymous):

ehh, im not good with trig :/

OpenStudy (amistre64):

the trig itself isnt the issue; they want you to become familiar with identites; this really just reads as this: ab^2 - a^3 = a ; when b^2 = 1-a^2 a(1-a^2) - a^3 = a a -a^3 + a^3 = a a = a

OpenStudy (amistre64):

cos^2 = 1 - sin^2

OpenStudy (anonymous):

mkay, so far i get that..

OpenStudy (amistre64):

sin*cos^2 + sin^3 = sin ; and cos^2 = 1-sin^2 sin*(1-sin^2) + sin^3 = sin sin -sin^3 + sin^3 = sin sin = sin helps if i dont typo it

OpenStudy (anonymous):

thank you so much!

OpenStudy (amistre64):

yep